OpenStudy (anonymous):

HELPPPPPP ANYONE !!! Each statement describes a transformation of the graph of y = x. Which statement correctly describes the graph of y = x - 16? It is the graph of y = x translated to the right 16 units. It is the graph of y = x translated up 16 units. It is the graph of y = x, where the slope is decreased by 16. It is the graph of y = x translated to the left 16 units

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

\(\large\color{ slate }{\large {\bbox[5pt, ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= bx ~~~~\rm{\Rightarrow}~~~~ g(x)= b(x \normalsize\color{red }{ -~\rm{c} }) &~\rm{to~~the~~right~} \\ \text{ } \\f(x)= bx ~~~~\rm{\Rightarrow}~~~~ g(x)= b(x \normalsize\color{red }{ +~\rm{c}) } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= bx ~~~~\rm{\Rightarrow}~~~~ g(x)= bx \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= bx ~~~~\rm{\Rightarrow}~~~~ g(x)= bx \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

this is a rule for shifting a line f(x)=bx any direction. (where b is the coefficient of x).

OpenStudy (solomonzelman):

in your case this coefficient is 1. And this, your shift left from \(\large\color{slate}{ y=1x }\) would be \(\large\color{slate}{ y=1(x+c) }\) and the shift up would be \(\large\color{slate}{ y=1x+c }\) bt you know that \(\large\color{slate}{ y=1x+c }\) is same as \(\large\color{slate}{ y=1(x+c) }\) as you expand the parenthesis.

OpenStudy (solomonzelman):

And same way it happens when you have a shift light and shift down in this case, are same. to the right by c units. \(\large\color{slate}{ y=1(x-c ) }\) down by c units \(\large\color{slate}{ y=1x-c }\) As you expand the parenthesis, \(\large\color{slate}{ y=1(x-c) }\) is same as \(\large\color{slate}{ y=1x-c }\) and